# On Topology and Geometry of Hessian Manifolds

### v1

@@ 10 Riemannian Geometric Aspect
Let $((M,\nabla),h)$ be a Hessian manifold, and let $A$ be its Amari–Chentsov tensor. Then $A$ is totally symmetric and
@@
$K=-\frac{\left|\begin{array}[]{ccc}f_{xx}&f_{xxx}&f_{xxy}\\ f_{xy}&f_{xxy}&f_{xyy}\\ f_{yy}&f_{xyy}&f_{yyy}\end{array}\right|}{4\left|\begin{array}[]{cc}f_{xx}&f_{xy}\\ f_{xy}&f_{yy}\end{array}\right|^{2}}.$
@@
Let $((M,\nabla),h)$ be a compact orientable Hessian manifold. Then $M$ is a mapping torus.
Corollary 10.10.
@@ 11 The Cheng-Yau Metric
For any $k>0$, define function $u_{k}\in C^{\infty}(\Omega)$ by $u_{k}(\textup{{x}}):=u(k\cdot\textup{{x}})+\operatorname{log}(k)$. Then straightforward computation yields that $u_{k}$ is also a Cheng-Yau potential of $\Omega$ for every $k>0$. Therefore, by the uniqueness part of Theorem11.1, we obtain the functional equation
$u(\textup{{x}})-u(k\cdot\textup{{x}})=\operatorname{log}(k)$
@@
$\frac{u(\textup{{x}}+k\cdot\textup{{x}})-u(\textup{{x}})}{k}=-\operatorname{log}(1+k)/k\rightarrow-1$
as $k\rightarrow 0$, we have that
@@
12 Compact Examples of Koszul Type
@@
The flat tori, or more generally, Bieberbach manifolds are examples of compact Hessian manifolds, but they are trivial in the sense that their affine structures are given by their Levi-Civita connections. In is section, we construct concretely a family of compact Hessian manifolds of Koszul type with negative scalar curvature.
@@
Recall from the theory of Lie groups that the isometry group $\operatorname{PSL}_{2}(\mathbb{R})$ of $\mathbb{H}^{2}$ is isomorphic to the Lorentz group $\operatorname{SO}^{+}(2,1)$, whilst the latter is a Lie subgroup of $\operatorname{GL}_{3}(\mathbb{R})$. It follows that $G$ acts on the convex cone $\Omega$ by linear automorphisms. Therefore, $\tilde{\nabla}$ descends to a flat and torsion-free connection $D$ on the tangent bundle of the mapping torus $M_{\varphi}=(\mathbb{H}^{2}\times\mathbb{R})/G$ of $\varphi$, so that $(M_{\varphi},D)$ is an affine manifold.
@@
This proves that $((M_{\varphi},D),h)$ is a compact Hessian manifold of Koszul type with constant negative scalar curvature 1-1, as claimed.
@@
It is obvious that the product of two Hessian manifolds is naturally a Hessian manifold. However, the converse statement is far from being correct, as in the example above, the product of the circle $\mathbb{S}^{1}$ and a closed orientable surface $\Sigma_{g}$ of genus $g>1$ admits an affine structure, but $\Sigma_{g}$ itself does not admit any affine structure if $g\neq 1$ by the proof of Chern’s conjecture in dimension two.
@@
13 Exotic Potentials for the Euclidean plane
@@
Consider the smooth mapping $\gamma:=\textbf{x}/\|\textbf{x}\|$. By the assumption on vectors $\textbf{x},\textbf{x}_{u},\textbf{x}_{v}$, a straightforward computation shows that $\gamma_{u}\times\gamma_{v}\equiv 0$. Therefore, the image $\gamma(\Omega)$ is a smooth curve on the $2$-sphere $\mathbb{S}^{2}:=\{(x,y,z)\in\mathbb{R}^{3}|x^{2}+y^{2}+z^{2}=1\}$. By the implicit function theorem, there exists a smooth function $f\colon\mathbb{S}^{2}\rightarrow\mathbb{R}$ such that $\gamma(\Omega)$ is locally the zero locus of $f$. Now, define homogeneous function $F$ of degree zero via
@@
is a Lorentzian metric. Moreover, since
$\frac{f_{xx}+f_{yy}}{\sqrt{f_{xx}f_{yy}-f_{xy}^{2}}}\equiv 0,$
by Proposition13.2, $g$ is flat, i.e. locally the metric of the $1+1$ Minkowski space.
@@
of $f$ in to the flat metric tensor $dr^{2}+r^{2}d\theta^{2}$ of the Euclidean plane in its polar coordinates. The flatness of $g$ can be proved also by noticing that
@@
14 Topology of Hessian 4-Manifolds
@@
The first claim follows from Proposition2.5&5.7 and Hirzebruch’s signature theorem. Moreover, since $\operatorname{dim}M=4$ is even and $\chi(M)=0$, by Poincaré duality the first Betti number of $M$ is positive and the second Betti number of $M$ is even. ∎
@@
Let $((M,\nabla),h)$ be an oriented closed Hessian 4-manifold. If $M$ is spin, then it holds that $Q_{M}=H^{\oplus m}$ where
$H:=Q_{\mathbb{S}^{2}\times\mathbb{S}^{2}}=\left[\begin{array}[]{cc}0&1\\ 1&0\end{array}\right]$
and $2m$ is the second Betti number of $M$. If $M$ is not spin, then
$Q_{M}=\operatorname{diag}(\underbrace{1,\dots,1}_{m},\underbrace{-1,\dots,-1}_{m})$
where $2m$ is the the second Betti number of $M$.
@@
Let $((M,\nabla),h)$ be a closed orientable Hessian-Einstein 4-manifold. Then $(M,h)$ is a flat Riemannian manifold, and $\nabla$ is its Levi-Civita connection.
@@
Assume the contrary that the Einstein manifold $(M,h)$ is not flat. Then, since by Lemma14.1

### v2

@@ 3.5 Riemannian Geometric Aspect
Let $((M,\nabla),h)$ be a Hessian manifold, and $A$ its Amari–Chentsov tensor. Then $A$ is totally symmetric and
@@
$K=-\frac{\left|\begin{array}[]{ccc}f_{xx}&f_{xxx}&f_{xxy}\\ f_{xy}&f_{xxy}&f_{xyy}\\ f_{yy}&f_{xyy}&f_{yyy}\end{array}\right|}{4\left|\begin{array}[]{cc}f_{xx}&f_{xy}\\ f_{xy}&f_{yy}\end{array}\right|^{2}}=\frac{f_{xx}\{f_{yy},f_{xy}\}+f_{yy}\{f_{xy},f_{xx}\}+f_{xy}\{f_{xx},f_{yy}\}}{4(f_{xx}f_{yy}-f_{xy}f_{xy})^{2}}$
where $\{\cdot,\cdot\}$ is the standard Poisson bracket normalized by $\{x,y\}=1$.
@@ 3.6 The Cheng-Yau Metric
For any $t>0$, define function $u_{t}\in C^{\infty}(\Omega)$ by $u_{t}(\textup{{x}}):=u(t\cdot\textup{{x}})+\operatorname{log}(t)$. Then straightforward computation yields that $u_{t}$ is also a Cheng-Yau potential of $\Omega$ for every $k>0$. Therefore, by the uniqueness part of Theorem3.43, we obtain the functional equation
$u(\textup{{x}})-u(t\cdot\textup{{x}})=\operatorname{log}(t)$
@@
$\frac{u(\textup{{x}}+t\cdot\textup{{x}})-u(\textup{{x}})}{t}=-\operatorname{log}(1+t)/t\rightarrow-1$
as $t\rightarrow 0$, we have that
@@
Using Lemma3.45, we can prove that Cheng-Yau potentials are harmonic functions.
Proposition 3.46.
Let $\Omega$ be a domain in $\mathbb{R}^{d}$, and let $u\in C^{\infty}(\Omega)$. Suppose that $\operatorname{Hess}(u)$ is nowhere degenerated, so that $g=(\partial_{i}\partial_{j}u)dx^{i}\otimes dx^{j}$ is a pseudo-Riemannian metric on $\Omega$. Then, the Laplace–Beltrami operator $\Delta^{g}$ of $(\Omega,g)$ satisfies
$\Delta^{g}\phi=\operatorname{tr}(\operatorname{Hess}(u)^{-1}\operatorname{Hess}(\phi))-\frac{1}{2}\langle\nabla\operatorname{log}(\operatorname{det}(\operatorname{Hess}(u))),\operatorname{Hess}(u)^{-1}\nabla\phi\rangle$
for all $\phi\in C^{\infty}(\Omega)$.
Proof.
Proposition 3.47.
Let $\Omega$ be a regular convex cone in $\mathbb{R}^{d}$, and let $u\in C^{\infty}(\Omega)$ be the Cheng-Yau potential of $\Omega$, so that $g=(\partial_{i}\partial_{j}u)dx^{i}\otimes dx^{j}$ is a Riemannian metric on $\Omega$. Then $u$ is a harmonic function on $(\Omega,g)$.
Proof.
Since $u$ is a Cheng-Yau potential, we have $2d\cdot u=\operatorname{log}(\operatorname{det}(\operatorname{Hess}(u)))$. Denote by $\Delta^{g}$ the Laplace–Beltrami operator of $(\Omega,g)$. Then by Lemma3.45 and Proposition3.46,
$\Delta^{g}u=$ $\operatorname{tr}(\operatorname{Hess}(u)^{-1}\operatorname{Hess}(u))-\frac{1}{2}\langle\nabla\operatorname{log}(\operatorname{det}(\operatorname{Hess}(u))),\operatorname{Hess}(u)^{-1}\nabla u\rangle$
$=$ $\operatorname{tr}(\operatorname{Id})-d\cdot\langle\nabla u,\operatorname{Hess}(u)^{-1}\nabla u\rangle$
$=$ $d-d\cdot 1$
$=$ $0$
as required. ∎
@@
4 Topics in Low Dimensional Hessian Geometry
4.1 Complete Hessian surfaces
It is a well-known fact that an orientable closed Hessian surface is homeomorphic to the 2-torus $\mathbb{T}^{2}$. The ultimate objective of this section is to classify all complete Hessian surfaces topologically. To achieve this goal, the following preparations are needed:
Definition 4.1.
Let $(M,\nabla)$ be an affine manifold. We denote by $\operatorname{Aff}(M;\nabla)$ the group of affine automorphisms of $(M,\nabla)$. Whenever $M=\Omega$ is a convex domain in $\mathbb{R}^{d}$, and $\nabla$ is the Levi-Civita connection of the $d$-dimensional flat Euclidean space, by convention we omit $\nabla$ to write $\operatorname{Aff}(\Omega):=\operatorname{Aff}(\Omega;\nabla)$.
Lemma 4.2.
Let $\Omega$ be a convex domain in $\mathbb{R}^{d}$ and $\operatorname{recc}(\Omega)$ its recession cone. Let $f\in\operatorname{Aff}_{2}(\mathbb{R})$ and let $T$ be the linear part of $f$, that is, $f(\textup{{x}})=T\textup{{x}}+f(0)$ for all $\textup{{x}}\in\mathbb{R}^{d}$. Suppose that $\Omega=f(\Omega)$. Then $T\textup{{x}}\in\operatorname{recc}(\Omega)$ for all $\textup{{x}}\in\operatorname{recc}(\Omega)$.
Proof.
Take any $\textbf{x}\in\operatorname{recc}(\Omega)$ and $\textbf{y}\in\Omega$. By the definition of a recession cone, it suffices to prove that $T\textbf{x}+\textbf{y}\in\Omega$. But indeed, we have
$T\textbf{x}+\textbf{y}=T\textbf{x}+Tf^{-1}(\textbf{y})+f(0)=f(\textbf{x}+f^{-1}(\textbf{y}))\in\Omega$
as required. ∎
The following proposition concerning compactness of the affine automorphism group of a bounded convex domain is itself interesting and useful.
Proposition 4.3.
Let $\Omega$ be a bounded convex domain in $\mathbb{R}^{n}$. Then $\operatorname{Aff}(\Omega)$ is isomorphic to a compact Lie subgroup of $\operatorname{GL}_{n}(\mathbb{R})$.
Proof.
It is readily seen that $G:=\operatorname{Aff}(\Omega)$ is a closed subgroup of $\operatorname{Aff}_{n}(\mathbb{R})$. By Cartan’s theorem on closed subgroups, $G$ is a Lie subgroup of $\operatorname{Aff}_{n}(\mathbb{R})$.
Denote by $\overline{\Omega}$ the closure of $\Omega$. By Heine-Borel theorem and the assumption on $\Omega$, we have that $\overline{\Omega}$ is a convex body. Denote by $E_{\Omega}$ the inner Löwner–John ellipsoid of $\overline{\Omega}$. By the uniqueness of the inner Löwner–John ellipsoid, $E_{\Omega}=g(E_{\Omega})$ for every $g\in G$. By linear algebra, if a non-degenerate affine transform $g$ maps an ellipsoid $E$ onto $E$ itself, then $g$ fix the center of $E$. It follows that all elements of $G$ share a common fixed point, namely the center of the ellipsoid $E_{\Omega}$. Without lose of generality, we may assume that the ellipsoid $E_{\Omega}$ is centered at the origin $0\in\mathbb{R}^{n}$. Therefore $G$ is a subgroup of $\operatorname{GL}_{n}(\mathbb{R})$.
Denote by $\|\cdot\|_{\operatorname{op}}$ the operator norm. Denote by $\ell_{1}$ and $\ell_{2}$ the diameter and the width of $E_{\Omega}$, respectively. Then it is readily seen that $\|g\|_{\operatorname{op}}\leq\ell_{1}/\ell_{2}$ for all $g\in G$. In particular, $G$ is a bounded subset of $\mathbb{R}^{n\times n}$. Now, again by Heine-Borel theorem, $G$ is compact. ∎
We will also use the following classical theorem on crystallographic groups of D. Fried and W. Goldman.
Theorem 4.4 (Fried-Goldman,1983[28]).
Let $G$ be a subgroup of $\operatorname{Aff}(\mathbb{R}^{2})$. Suppose that $G$ acts proper discontinuously on $\mathbb{R}^{2}$ by affine transform. Then $G$ admits a solvable subgroup of finite index.
Proof.
This is a main theorem proved in [28]. ∎
Now we are fully prepared to prove the main theorem of this section.
Theorem 4.5.
Let $((M,\nabla),h)$ be a complete Hessian surface. Then $M$ is diffeomorphic to one of the following smooth manifolds:
  1. 1.
    the plane $\mathbb{R}^{2}$;
  2. 2.
    the cylinder $\mathbb{S}^{1}\times\mathbb{R}$;
  3. 3.
    the Möbius strip;
  4. 4.
    the 2-torus $\mathbb{T}^{2}$;
  5. 5.
    the Klein bottle.
Proof.
Since $((M,\nabla),h)$ is a complete Hessian manifold of dimension 2, by Theorem2.13, there exists a planar convex domain $\Omega$ and a discrete subgroup $G$ of the affine group $\operatorname{Aff}_{2}(\mathbb{R})$, such that $G$ acts freely and proper discontinuously on $\Omega$ by affine transform, and that $M\cong\Omega/G$.
If $G$ is trivial, then $M=\Omega$ is homeomorphic to $\mathbb{R}^{2}$. If $M$ is closed, then by Theorem2.17, we have that $\chi(M)=0$ and hence $M$ is homeomorphic to the 2-torus or the Klein bottle. Recall from low-dimensional topology that the fundamental group of a non-compact surface is always free. From now on, we assume that $M$ is non-compact and $G$ is a non-trivial free group.
Denote by $\operatorname{recc}(\Omega)$ the recession cone of $\Omega$. By elementary Euclidean geometry, any convex cone in $\mathbb{R}^{2}$, and hence in particular $\operatorname{recc}(\Omega)$, is congruent to one of the following sets:
  1. 1.
    a point in $\mathbb{R}^{2}$;
  2. 2.
    a straight line, i.e. an 1-dimensional affine subspace of $\mathbb{R}^{2}$;
  3. 3.
    a wedge, i.e. the set $W_{\theta}:=\{z\in\mathbb{C}^{1}:0\leq\operatorname{Arg}(z)\leq\theta\}$ for $0\leq\theta\leq\pi$;
  4. 4.
    the whole plane $\mathbb{R}^{2}$.
Assume that $\operatorname{recc}(\Omega)$ is a point. Then, by convex geometry, we have that $\Omega$ is bounded. Therefore, by Proposition4.3, $\operatorname{Aff}(\Omega)$ is a compact Lie group. But, by assumption, $G$ is a infinite discrete subgroup of $\operatorname{Aff}(\Omega)$, which leads to a contradiction. Therefore $\operatorname{recc}(\Omega)$ cannot be a point.
Assume that $\operatorname{recc}(\Omega)$ is the whole plane $\mathbb{R}^{2}$. Then we have $\Omega=\operatorname{recc}(\Omega)=\mathbb{R}^{2}$. Therefore $G$ admits a solvable subgroup $G_{0}$ of finite index by Theorem4.4.
Assume that $\operatorname{recc}(\Omega)$ is a straight line in $\mathbb{R}^{2}$ or a wedge $W_{\theta}$ with $0\leq\theta\leq\pi$. Consider the group homomorphism
$\phi\colon\operatorname{Aff}_{2}(\mathbb{R})\rightarrow\operatorname{GL}_{2}(\mathbb{R}),\quad f\mapsto(\textbf{x}\mapsto f(\textbf{x})-f(0))$
that is, the map sending every affine transform to its linear part. Then by Lemma4.2, we have that $\phi(G)$ acts on $\operatorname{recc}(\Omega)$ by linear automorphism. In particular, $\phi(G)$ preserves the boundary of $\operatorname{recc}(\Omega)$. Therefore, $\phi(G)$ admits a subgroup $H$ of index at most 2, such that all elements of $H$ share at least one eigenvector, and hence $G_{0}:=\phi^{-1}(H)\cap G$ is isomorphic to a subgroup of the group $B$ of invertible upper-triangular real $3\times 3$ matrices. Recall from linear algebra that the length of the derived sequence of Heisenberg group $H_{3}(\mathbb{R})$ is exactly $2$, and observe that $B/H_{3}(\mathbb{R})\cong(\mathbb{R}^{\times})^{3}$ is Abelian. By group theory we obtain that $B$ is solvable. Since $B$ is solvable and $G_{0}$ is a subgroup of $G$ of index at most 2, we again have that $G$ admits a solvable subgroup $G_{0}$ of finite index.
This proves that $G$ always admits a solvable subgroup $G_{0}$ of finite index.
Since $G$ is infinite, we have $G_{0}$ ditto. By Nielsen–Schreier theorem, $G_{0}$ is also free. Since any free group of rank $r>1$ is not solvable, the rank of $G_{0}$ is at most 1. Therefore, we conclude that $G_{0}\cong\mathbb{Z}$. Finally, by Schreier’s index formula, we arrive at that in fact the rank of $G$ is also 1, i.e. $G$ is isomorphic to $\mathbb{Z}$.
Therefore, the non-compact surface $M$ has fundamental group $\pi_{1}(M)=\mathbb{Z}$.
By the classification theorem of non-compact surfaces, we have that $M$ is a closed surface $N$ with $k\geq 1$ punctures.
Assume that $N$ is orientable of genus $g$. Then by Seifert-van Kampen theorem, $\pi_{1}(M)$ is a free group of rank r=2g+k1r=2g+k-1. Since $r=1$ and $k>0$, we have that $g=0$ and $k=2$. Therefore $M$ is a sphere with $2$ punctures, i.e. a cylinder.
Assume that $N$ is non-orientable. Then, by the classification theorem of closed surfaces, $N$ is the connected sum of $l$ real projective planes for some integer $l\geq 1$. Again by Seifert-van Kampen theorem, $\pi_{1}(M)$ is a free group of rank r=l+k1r=l+k-1. Since $r=1$ and $k>0$, we obtain $l=k=1$. Therefore $M$ is a punctured $\mathbb{R}P^{2}$, i.e. a Möbius strip.
Quod erat demonstrandum. ∎
Example 4.6.
A complete Hessian surface of constant Gaußian curvature 1-1 can be isometric to a hyperbolic funnel, that is, the quotient of the Lobachevsky plane by an expanding homothety $z\mapsto\lambda\cdot z$ where $\lambda>1$.
In the polar coordinates $(r,\theta)$ of the complex plane, the $1$-form potential of a hyperbolic funnel is precisely
$\eta=\frac{dr}{2r}+\left(\frac{1}{2}\frac{\operatorname{sin}(\theta)}{\operatorname{cos}(\theta)}-\frac{\operatorname{sin}(\theta)}{\operatorname{cos}(\theta)^{3}}\right)d\theta$
defined on an open annulus centered at the origin. A straightforward computation shows that $\eta$ is closed, but not exact.
4.2 Compact Radiant Examples of Koszul Type
@@
The flat tori, or more generally, Bieberbach manifolds are examples of compact Hessian manifolds, but they are trivial in the sense that their affine structures are given by their Levi-Civita connections. In is section, we construct concretely a family of compact radiant Hessian manifolds of Koszul type with negative scalar curvature.
@@
Recall from the theory of Lie groups that the isometry group $\operatorname{PSL}_{2}(\mathbb{R})$ of $\mathbb{H}^{2}$ is isomorphic to the Lorentz group $\operatorname{SO}^{+}(2,1)$, whilst the latter is a Lie subgroup of $\operatorname{GL}_{3}(\mathbb{R})$. It follows that $G$ acts on the convex cone $\Omega$ by linear automorphisms. Therefore, $\tilde{\nabla}$ descends to a flat and torsion-free connection $D$ on the tangent bundle of the mapping torus $M_{\varphi}=(\mathbb{H}^{2}\times\mathbb{R})/G$ of $\varphi$, so that $(M_{\varphi},D)$ is an affine manifold. Moreover, $(M_{\varphi},D)$ is radiant by Proposition3.10.
@@
This proves that $((M_{\varphi},D),h)$ is a compact radiant Hessian manifold of Koszul type with constant negative scalar curvature 1-1, as claimed.
@@
It is obvious that the product of two Hessian manifolds is naturally a Hessian manifold. However, the converse statement is far from being correct, as in the example above, the product of the circle $\mathbb{S}^{1}$ and a closed orientable surface $\Sigma_{g}$ of genus $g>1$ admits an affine structure, but $\Sigma_{g}$ itself does not admit any affine structure if $g\neq 1$ by the proof of Chern’s conjecture in dimension two [20].
@@
4.3 Exotic Potentials for the Euclidean plane
@@
Consider the smooth mapping $\phi:=\textbf{x}/\|\textbf{x}\|$. By the assumption on vectors $\textbf{x},\textbf{x}_{u},\textbf{x}_{v}$, a straightforward computation shows that $\phi_{u}\times\phi_{v}\equiv 0$. Therefore, the image $\phi(\Omega)$ is a smooth curve on the $2$-sphere $\mathbb{S}^{2}:=\{(x,y,z)\in\mathbb{R}^{3}|x^{2}+y^{2}+z^{2}=1\}$. By the implicit function theorem, there exists a smooth function $f\colon\mathbb{S}^{2}\rightarrow\mathbb{R}$ such that $\phi(\Omega)$ is locally the zero locus of $f$. Now, define homogeneous function $F$ of degree zero via
@@
is a Lorentzian metric. Moreover, by Proposition4.10, $g$ is flat, i.e. locally the metric of the $1+1$ Minkowski space.
@@
of $f$ to the flat metric tensor $dr^{2}+r^{2}d\theta^{2}$ of the Euclidean plane in its polar coordinates. The flatness of $g$ can be proved also by noticing that
@@
4.4 Topology of Hessian 4-Manifolds
@@
The first claim follows from Proposition1.5&2.17 and Hirzebruch’s signature theorem. Moreover, since $\operatorname{dim}M=4$ is even and $\chi(M)=0$, by Poincaré duality the first Betti number of $M$ is positive and the second Betti number of $M$ is even. ∎
@@
Let $((M,\nabla),h)$ be an oriented closed Hessian 4-manifold with the second Betti number $b_{2}(M)=2m$. If $M$ is spin, then it holds that $Q_{M}=H^{\oplus m}$ where
$H:=Q_{\mathbb{S}^{2}\times\mathbb{S}^{2}}=\left[\begin{array}[]{cc}0&1\\ 1&0\end{array}\right];$
if $M$ is not spin, then $Q_{M}=\operatorname{diag}(\underbrace{1,\dots,1}_{m},\underbrace{-1,\dots,-1}_{m})$.
@@
Let $((M,\nabla),h)$ be a closed Hessian-Einstein manifold of dimension four. Then $(M,h)$ is a flat Riemannian manifold, and $\nabla$ is its Levi-Civita connection.
@@
By considering the orientation double cover, we may assume without lose of generality that $M$ is orientable.
Now, assume the contrary that the Einstein manifold $(M,h)$ is not flat. Then, since by Lemma4.15